OpenStudy (anonymous):

\[\left| 6x+2 \right|-2=6\]

OpenStudy (zepp):

Try to isolate the |6x + 2| by taking out -2 So it would be |6x + 2| = 6 +2 |6x + 2| = 8 We can conclude that 6x + 2 has two answers, 8 and -8 6x + 2 = 8; 6x + 2 = -8 6x + 2 = 8 6x = 8-2 6x = 6 x = 1 6x + 2 = -8 6x = -8 -2 6x = -10 x = -10/6 x = -5/3 x = 1; -5/3